find the following

  • Input :220V AC
  • Output: 5V DC
  • Minimum   RL   = 300
  • Ripple factor   r = 0.05

Measured Calculated Vecout) max Vout) min Vr PP Average Voltage (VDC)
Answer #1

First, let's start our discussion with some basics stuff before directly going to the solution.

ARC Ripple voltage (Van) = UPT Y = (ba-avy) 1 2x300x2+2x103 Voi = 0.130 (if Vj = 074) Dav 22-94%E ( جلال

I hope you understand the whole discussion and my handwriting. I hope this will be useful to you.

Thanking You